What amount of H2O is present when it takes 8328 joules to boil water to steam at 100 °C
1
Expert's answer
2020-06-02T14:08:57-0400

Heat of vaporization of water "H_v = 2260 \frac{J}{g}"

"q=m \times H_v"

"m=\frac{q}{H_v}= \frac{ 8328 J}{2260 \frac{J}{g}}=3.685 g"